OpenStudy (anonymous):

ALGEBRA 2 HELP!!! What is the discontinuity and zero of the function f(x) = the quantity of 3 x squared plus x minus 4, all over x minus 1? Discontinuity at (-1, 1), zero at (four thirds, 0) Discontinuity at (-1, 1), zero at (negative four thirds , 0) Discontinuity at (1, 7), zero at (four thirds , 0) Discontinuity at (1, 7), zero at (negtive four thirds, 0)

OpenStudy (anonymous):

all of the above

OpenStudy (anonymous):

??

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

\[\Large\rm f(x)=\frac{3x^2+x-4}{x-1}\]This is our function?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

To find both `zeroes` and `discontinuities` we set our function equal to zero,\[\Large\rm 0=\frac{3x^2+x-4}{x-1}\]We'll look at the numerator and denominator separately. The `numerator` will give us our zeroes. The `denominator` will give us our discontinuities. \[\Large\rm zeroes:\]\[\Large\rm 0=3x^2+x-4\]We'll have to throw this into the Quadratic Formula.\[\Large\rm x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]Remember how to do that?

OpenStudy (anonymous):

not really

zepdrix (zepdrix):

Our numerator is in the form:\[\Large\rm 0=ax^2+bx+c\]\[\Large\rm 0=3x^2+(1)x+(-4)\]

zepdrix (zepdrix):

So we'll plug all those values in, and try to simplify.

zepdrix (zepdrix):

\[\Large\rm x=\frac{-(1)\pm\sqrt{(1)^2-4(3)(-4)}}{2(3)}\]Understand where I got those numbers from? :o

OpenStudy (anonymous):

yeah

zepdrix (zepdrix):

So then uhhhh, what does that simplify down to? Do it!! c:

OpenStudy (anonymous):

i got (-1) i square root of -47 over 6 @zepdrix

OpenStudy (anonymous):

im pretty sure thats wrong

zepdrix (zepdrix):

Hmm under the root, see how we have subtraction AND a negative sign?

OpenStudy (anonymous):

yeah

zepdrix (zepdrix):

So that should change to positive, yes? :d

zepdrix (zepdrix):

\[\Large\rm x=\frac{-1\pm\sqrt{(1)^2+4(3)(4)}}{6}\]

OpenStudy (anonymous):

so -1 plus or minus 7 over 6

zepdrix (zepdrix):

\[\Large\rm x=\frac{-1+7}{6}\]\[\Large\rm x=\frac{-1-7}{6}\]Ok good, so that gives us two values, yes?

OpenStudy (anonymous):

so 1 and -8 over 6

zepdrix (zepdrix):

Ok great. And -8 over 6 will simplify to -4 over 3.

zepdrix (zepdrix):

We have to be careful though. I probably should have mentioned this earlier. It makes more sense to look for `discontinuities` FIRST. If we ever come up with the same value for a `discontinuity` and a `zero`, then we do not include that value as one of our zeroes.

zepdrix (zepdrix):

Let's jump to the denominator and see what's going on,\[\Large\rm 0=x-1\]

zepdrix (zepdrix):

What value do you get for your discontinuity?

OpenStudy (anonymous):

0

zepdrix (zepdrix):

I set the denominator equal to zero, now we solve for x.

OpenStudy (anonymous):

so we plug in what we just got?

zepdrix (zepdrix):

No. This is our denominator set equal to zero.\[\Large\rm 0=x-1\]Solve for x. Isolate the x. Get x alone.

OpenStudy (anonymous):

1

zepdrix (zepdrix):

Ok great, we find that our discontinuity is at x=1.

zepdrix (zepdrix):

But wait that was one of our zeroes! x=1, x=-4/3 That means we cannot include x=1 as a zero of the function.

zepdrix (zepdrix):

A discontinuity should NOT HAVE a y coordinate. Whoever wrote this question is a silly billy. We have a discontinuity at x=1. And a zero located at (-4,3, 0).

OpenStudy (anonymous):

so the answer is b!

zepdrix (zepdrix):

Woops I made a typo, lemme do that again :(

zepdrix (zepdrix):

And a zero located at (-4/3, 0).*

zepdrix (zepdrix):

Ummmmm

zepdrix (zepdrix):

Looks like it's D. See how the x coordinate is 1 for the discontinuity?

OpenStudy (anonymous):

yes thank you so much!

zepdrix (zepdrix):

cool c: